# Bridging Deep Reinforcement Learning and Motion Planning for Model-Free

### v1

@@ Abstract
Deep Reinforcement Learning (DRL) has emerged as a powerful model-free paradigm for learning optimal policies. However, in real-world navigation tasks, DRL methods often suffer from insufficient exploration, particularly in cluttered environments with sparse rewards or complex dynamics under system disturbances. To address this challenge, we bridge general graph-based motion planning with DRL, enabling agents to explore cluttered spaces more effectively and achieve desired navigation performance. Specifically, we design a dense reward function grounded in a graph structure that spans the entire state space. This graph provides rich guidance, steering the agent toward optimal strategies. We validate our approach in challenging environments, demonstrating substantial improvements in exploration efficiency and task success rates. The project website is available at: https://plen1lune.github.io/overcome_exploration/
Index Terms:
@@
Deep Reinforcement Learning, Motion Planning, Sampling-based Method
I Introduction
Recently, the demand for learning control policies increasingly prevalent across a wide range of real-world tasks. Examples include agile drone flight in challenging scenes [1], mobile robot navigation [2, 3, 4], and object manipulation [5]. These tasks often involve high-dimensional sensory inputs and uncertainty in dynamics, where classical control methods are insufficient. For such tasks, DRL has become a dominant framework for obtaining control policies, thanks to its model-free nature that allows agents to learn directly from interactions without relying on explicit dynamics models [6, 7]. This flexibility makes DRL particularly appealing in domains such as robotic manipulation [8, 9], autonomous driving [10], and locomotion [11], where constructing accurate models is often infeasible. In these applications, DRL offers the potential to generalize across tasks and adapt to uncertainties, establishing itself as a key approach for intelligent decision-making under complex conditions.
@@ I Introduction
However, the practical success of DRL remains severely limited by its sample inefficiency, primarily due to inadequate exploration in sparse-reward or high-dimensional settings [12, 13]. Without informative feedback, agents may spend substantial time interacting with irrelevant states. To address this issue, a range of exploration strategies have been proposed. Intrinsic motivation methods reward agents for visiting novel or uncertain states [14, 15], while count-based approaches approximate visitation frequencies using density models [16, 13]. Noise-driven methods, such as entropy regularization [17] or stochastic policy perturbation [18], also remain popular due to simplicity. Some studies decompose the exploration process either by partitioning the task into subtasks [19] or employ temporal logic (TL) to guide exploration [20].While these methods partially enhance sample efficiency, they fall short of resolving the fundamental difficulty of agent exploration in cluttered environments.
Classical motion planning algorithms, such as RRT, provide structured strategies for navigating cluttered environments by exploiting geometric information and long-horizon planning capabilities [21, 22]. However, these approaches are typically integrated with model-based path-tracking control approaches such as Model Predictive Control (MPC) [23]. Model-free path-tracking, which is the focus of this paper, is still an open problem. To address this limitation, Cai et al. [24] proposed a motion planning-guided DRL framework, where waypoints are generated via Rapidly-exploring Random Trees (RRT) to guide agent exploration. This integration provides structured, long-horizon guidance, boosting sample efficiency and enabling faster convergence, especially in maze-like or constrained environments [25]. The agent leverages both classical planning coverage and model-free policy adaptation to explore efficiently. However, the planned trajectories may themselves be infeasible due to environmental constraints, or become invalid under uncertainty [3, 26, 27]. This static nature of the path limits adaptability and robustness, and calls for learning frameworks that can handle dynamic replanning and online corrections. However, in certain real-world tasks, real-time replanning may be impractical due to computational constraints, latency requirements, or limited sensing capabilities, making it essential to develop exploration strategies that are both efficient and lightweight [28, 29, 30].
Contributions: In this work, we propose a novel graph-based framework that enhances RL exploration in cluttered environments and can be seamlessly integrated with a wide range of model-free RL algorithms. Compared to previous work [24], our method achieves more comprehensive exploration of the environment’s state space, allowing agents to better handle challenging scenarios and potential disturbances, while fully harnessing the capabilities of model-free reinforcement learning to learn optimal policies under unknown dynamics. We further provide theoretical guarantee that our exploration strategy preserves the original RL objective and accelerates training convergence. In addition, our framework enables agents to generalize across arbitrary initial states without retraining or policy modification.
Organization: Sec. II mentions the foundations of our approach. In Sec. III, we define the problem and emphasize the challenges. Sec. IV presents our approach for addressing simple goal-reaching tasks via reward design using sampling-based methods over the workspace. Sec. V demonstrate how we generalize the initialization positions. The performance of the proposed method is shown in Sec. VI.
II Related Works
Motion Planning: Sampling-based motion planning methods, such as Rapidly-exploring Random Tree (RRT) [31], RRT* [22], Probabilistic Road Map (PRM) [32], and Rapidly-exploring Random Graph (RRG) [22], are widely used for finding collision-free paths in continuous geometric spaces. Among these, PRM constructs a roadmap by sampling random points in the configuration space, connecting them through feasible edges based on a local planner, and using this graph representation to find paths between specified start and goal points. This method is particularly efficient for multi-query settings, as the precomputed roadmap can be reused across different planning tasks. RRG extends the idea of RRT by maintaining a graph structure that connects all sampled points, thereby ensuring asymptotic optimality while allowing more flexible pathfinding. To take advantage of these graph structures for navigation, shortest-path algorithms such as A* [33] are often employed to compute optimal paths on the graph. A* performs heuristic-guided search to efficiently identify paths that minimize cost, making it a powerful tool for motion planning tasks involving graph-based representations. While these methods provide robust solutions for high-dimensional motion planning, they typically rely on model-based path-tracking controllers to execute the planned trajectories.
@@ II Related Works
Deep Reinforcement Learning: Deep Reinforcement Learning (DRL) has emerged as a powerful model-free paradigm for learning optimal policies in complex decision-making problems [7]. In continuous control tasks, DRL algorithms such as Proximal Policy Optimization (PPO) [18] and Soft Actor-Critic (SAC) [17] have been widely adopted due to their empirical stability and sample efficiency. These algorithms are typically built upon the actor-critic framework, which has become the standard paradigm in modern policy-based reinforcement learning [7, 17]. In this framework, the actor is a parameterized policy that maps states to actions, while the critic commonly estimates the expected return—either in the form of a state-value function or an action-value function—to guide the improvement of the actor. This separation enables the actor to focus on policy optimization while leveraging the critic to provide accurate feedback based on the agent’s experience, thereby facilitating more effective and stable learning dynamics.
III Problem Formulation
@@
System Disturbance: We consider a continuous-time dynamical system $S$ with general unknown dynamics, whose evolution is described by:
$\frac{ds}{dt}=\dot{s}=f(s,a,d)$ (1)
where $s\in\mathcal{S}\subseteq\mathbb{R}^{n}$ represents the state variables, $a\in\mathcal{A}\subseteq\mathbb{R}^{m}$ denotes the control inputs, and $d$ is the disturbance function.
@@
Figure 1: Comparison between previous method and graph-based method. We implemented both RRT (left) and RRG (right) to solve the same goal-reach task. As shown, the gray areas represent known obstacles, and the red cross indicate potential disturbances that may lead to failure transitions.
@@
Consider a navigation task in a cluttered environment with unknown system dynamics. The goal is to learn the optimal controllers $\pi^{*}$ that enable safe navigation of goal-reaching tasks.
This problem concerns finding a feasible solution in challenging environments. While standard MDP theory guarantees the existence of an optimal policy under mild conditions, learning such a policy via RL remains difficult in sparse-reward or long-horizon settings. These challenges are further exacerbated in practice by the clutter of the environment. Our work specifically addresses these issues by enhancing exploration through a graph-based guidance framework.
@@
We compared our approach with methods that use motion planning-generated paths to guide the training of reinforcement learning agents [24] (see Figure 1). To illustrate the limitations, we deliberately introduced an infeasibility (indicated by a red cross) along the planned geometric path to the goal. The path-based method (left) fails to manage such disturbances, as it relies on a single, fixed trajectory without considering alternative routes, resulting in failure. Consequently, using waypoints from this path as guidance also fails. In contrast, our approach employs a graph that covers the entire workspace (right), providing more robust guidance.
IV Overcoming Exploration
In Section IV, we introduce the graph-based reward and our design enabling arbitrary starting points.
IV-A Geometric Graph-based Motion planning Methods
@@
Given two states $s$ and $s^{\prime}$, the function $\text{Steer}(s,s^{\prime},\eta)$ returns a new state $s_{\text{new}}$ such that $s_{\text{new}}$ lies on the path connecting $s$ to $s^{\prime}$ and the distance $\|s-s_{\text{new}}\|_{2}\leq\eta$. The step size $\eta$ is a user-defined parameter that controls how far the path extends in a single step.
Sampling and connection
@@
The sampling function generates a new state $s_{\text{new}}$ in the free space, and the connection function attempts to connect $s_{\text{new}}$ to its nearest neighbors $\mathcal{N}(s_{\text{new}})$ using a geometric trajectory. For RRG, if a connection is successful, the edge is added to the graph, forming a cycle when possible.
Given the graph $G=(V,E)$ constructed by RRG, if a node $s\in V$ lies within the goal region $\mathcal{S}_{G}$, we compute the optimal state trajectory $\mathbf{x}^{*}$ as a sequence of geometric states:
$\mathbf{x}^{*}=\{s_{0},s_{1},\ldots,s_{d}\},$
where $s_{0}$ is the start state, $s_{d}\in\mathcal{S}_{G}$, and $s_{i}\in V$ for $i=1,\ldots,d$. In most planning pipelines, Dijkstra’s algorithm [34] is employed to extract the shortest waypoint sequence from the constructed graph.
@@
Beyond sampling-based graphs, grid-based graphs constitute another practical instantiation of our framework. In such cases, connectivity is predefined, and shortest paths are typically computed using algorithms such as A*. Additional details are provided in Appendix -A.
IV-B Motion Planning Guided Reward
@@
After employing a graph-based motion planning method, we obtained a set of waypoints that span the entire graph. Based on this, we use an optimal path algorithm to compute the distance from each node to the goal region. We denote the optimal trajectory by $\mathbf{x^{*}}$, and the optimal is defined according to the length of trajectories. $s|\mathbf{x^{*}}$ denotes a state of trajectory $\mathbf{x^{*}}$. We define the distance function $Dist:\mathcal{S}\times\mathcal{S}\times\mathcal{X}\rightarrow[0,\infty]$ to represent the Euclidean distance between two states along a given trajectory, and define the cost function $Cost:\mathcal{S}\times\mathcal{X}\rightarrow[0,\infty]$ to represent the distance between one state and the destination $s_{d}$. Hence, the distance between two states along the optimal trajectory can be expressed as $Dist(s_{1},s_{2}|\mathbf{x^{*}})=|Cost(s_{1}|\mathbf{x^{*}})-Cost(s_{2}|\mathbf{x^{*}})|$. Consequently, the graph $G$ is now extended to $G=(V,E,Cost)$, where we slightly abuse the notation, denoting $Cost$ as distance between the current state and the goal.
@@
The distance along the trajectory enables us to accurately quantify the actual path length, making it more applicable in cluttered environments compared to the Euclidean distance. Based on this distance, we design the motion-planning reward scheme to guide the RL agent. Given the difficulty of reaching an exact state in continuous space, we define the norm r-ball for loosen the criteria for determining whether the agent has reached specific states. Formally, it is defined by $Ball_{r}(s)=\left\{s^{\prime}\in\mathcal{S}\mid\|s-s^{\prime}\|_{2}\leq r\right\}$, where $r$ is the radius. The selection of the r-ball radius can influence the performance of the algorithm. For simplicity, we use the default $r=\frac{\eta}{2}$ based on the steering function in the following sections. We then define a progression function $D:\mathcal{S}\rightarrow[0,\infty]$ to determine whether the agent had moved toward the goal region and got closer to it. Formally,
@@
We adopt the definition of history from[35] with a slight modification $h_{t}=\{s_{1},a_{1},\dots,s_{t-1},a_{t-1}\}$ instead of $\{s_{1},a_{1},\dots,s_{t-1},a_{t-1},s_{t}\}$ , thereby the $D_{min}$ over a state-action sequence $\tau_{t}=h_{t}\bigcup\{s_{t},a_{t}\}$ is defined as $D_{min}(\mathbf{\tau})=\min_{s\in\mathbf{s}}{D(s)}$, where $\mathbf{s}$ denotes the state sequence in trajectory $\tau_{t}$. Ideally, we would check $D_{min}$ at each step to determine whether the agent has moved closer to the goal region, i.e., check if $D_{min}(\tau_{t})<D_{min}(\tau_{t-1})$, and assign a positive reward $R_{+}$ accordingly. However, this design introduces a non-Markovian reward issue. In other words, $r(h_{t},s_{t},a_{t})\neq r(h^{\prime}_{t},s_{t},a_{t})$ could happen with different history $h_{t}$ and $h^{\prime}_{t}$. To address this problem, we denote $\mathcal{S}^{\times}=\{(s,D_{min})\mid s\in\mathcal{S},\,D_{min}\in\mathbb{R}\}$, so that $D_{min}$ can be directly retrieved from the augmented state. Correspondingly, the augmented reward function is defined as $\tilde{r}:\mathcal{S}^{\times}\times\mathcal{A}\times\mathcal{S}^{\times}\to\mathbb{R}$,
@@
This result ensures that our planning-augmented reward remains compatible with standard RL algorithms. In the remainder of this paper, we operate in the augmented state space $\mathcal{S}^{\times}$ to maintain this property. The detailed proof is presented in Appendix -B.
@@
This proposition guarantees that the auxiliary rewards accelerate convergence without altering the original task objective. The proof is presented in Appendix -C. The learning objective aims to find the optimal policy by optimizing the parameters $\theta$:
@@
IV-C Infeasible Paths Analysis
Recalling Eq. (1), our method specifically addresses the challenge of unexpected disturbances in the simulation, which often render paths infeasible. In contrast to tree-based methods, which focus on planning a single path without accounting for its robustness or feasibility (e.g., with strong winds, as illustrated in Figure 4 (right) or other disturbances along the path), our approach leverages a graph-based motion planning strategy. By incorporating a comprehensive waypoint coverage across the entire map, the agent ensures that multiple potential paths are explored and evaluated during training.
We highlight that the robustness of our approach arises from the redundancy inherent in graph-based planning, which effectively mitigates the risk of path failure under disturbances. Instead of relying on a single predetermined trajectory, the agent benefits from multiple alternative paths towards the goal. As the density of waypoints and connectivity of the graph increase, the number of these alternative paths grows accordingly, greatly enhancing the likelihood that at least one safe and feasible path to the goal exists. This multiplicity ensures that even if certain paths become infeasible due to external perturbations, such as environmental disturbances, other viable paths remain available.

### v2

@@ Abstract
Deep Reinforcement Learning (DRL) has emerged as a powerful model-free paradigm for learning optimal policies. However, in navigation tasks with cluttered environments, DRL methods often suffer from insufficient exploration, especially under sparse rewards or complex dynamics with system disturbances. To address this challenge, we bridge general graph-based motion planning and DRL, enabling agents to explore cluttered spaces more effectively and achieve desired navigation performance. Specifically, we design a dense reward function grounded in a graph structure that spans the entire state space. This graph provides rich guidance, steering the agent toward optimal strategies. We validate our approach in challenging environments, demonstrating substantial improvements in exploration efficiency and task success rates.
@@
Keywords: Deep Reinforcement Learning, Motion Planning, Model-free Control
1 Introduction
Deep Reinforcement Learning (DRL) provides a flexible and general approach for acquiring control policies in tasks characterized by high-dimensional observations and complex, uncertain dynamics [1, 2]. Its model-free nature enables agents to learn directly from interactions, eliminating the need for explicit dynamics modeling. This makes DRL particularly suitable for domains such as robotic manipulation [3, 4], autonomous driving [5], and locomotion [6], where accurate models can be difficult to obtain. In these applications, DRL demonstrates strong potential for generalization and robustness, supporting intelligent decision-making under diverse and dynamic conditions.
@@ 1 Introduction
However, the practical success of DRL remains severely limited by its sample inefficiency, primarily due to inadequate exploration in sparse-reward or high-dimensional settings [7, 8]. Without informative feedback, agents may spend substantial time interacting with irrelevant states. Numerous strategies have been proposed to improve exploration in DRL, including intrinsic rewards [9, 10], density-based bonuses [11, 8], entropy regularization [12], policy perturbation [13], and hierarchical or logic-guided approaches [14, 15]. These approaches have partially alleviated exploration challenges and improved sample efficiency in regular environments. However, these methods still struggle with sparse reward signals and high-dimensional, cluttered environments—challenges that remain largely unsolved.
The primary contribution of this work is a graph-based exploration framework that enables efficient navigation from arbitrary initial states to task goals in cluttered environments. We address three key challenges that prior methods leave unresolved: 1) Bridging the gap between motion-planning feedback and the Markovian reward structures required by model-free RL algorithms; 2) handling infeasible planning paths and tracking errors; 3) enabling generalization to unseen cases. Our method builds on a classical motion planning backbone to provide informative guidance for exploration. It is lightweight, compatible with a range of model-free RL algorithms, and requires no policy modification across different initializations. We further provide theoretical justification that our exploration strategy preserves the original RL objective and empirically show improved sample efficiency and task success across a range of cluttered environments.
2 Related Works
Motion Planning: Sampling-based planners such as RRT, RRT*, PRM, and RRG [16, 17, 18] are widely used for collision-free path planning in continuous spaces. PRM builds a reusable roadmap for multi-query tasks by connecting sampled points with feasible edges, while RRG extends RRT with a graph structure for asymptotic optimality. Graph search algorithms like A* [19] are typically used to compute cost-optimal paths on these graphs. Although effective, these planners usually require model-based controllers to execute the planned trajectories.
@@ 2 Related Works
Deep Reinforcement Learning: DRL has emerged as a powerful model-free paradigm for learning optimal policies in complex decision-making problems [2]. In continuous control tasks, DRL algorithms such as Proximal Policy Optimization (PPO) [13] and Soft Actor-Critic (SAC) [12] have been widely adopted due to their empirical stability and sample efficiency. These algorithms are typically built upon the actor-critic framework, which has become the standard paradigm in modern policy-based reinforcement learning [2, 12]. In this framework, the actor is a parameterized policy that maps states to actions, while the critic commonly estimates the expected return—either in the form of a state-value function or an action-value function—to guide the improvement of the actor.
Exploration Strategies: A broad class of methods augment the reward signal to improve exploration. Curiosity-driven methods use forward dynamics models to measure prediction error as an intrinsic reward signal for novelty [9, 10, 20], encouraging agents to seek transitions that are poorly understood. Count-based exploration estimates state novelty through density modeling [11, 8], which generalizes tabular counts to continuous domains. Entropy-based methods [12] promote stochastic policies by maximizing uncertainty in action selection, while methods like PPO [13] introduce noise via clipped policy updates. Modular methods [14] introduce abstract goals or subtasks to guide long-horizon behavior, and diffusion-based methods [21] generate trajectories from good replay states to guide planning-based exploration. Balsells et al. [22] introduce the notion of reachability, whereas Gieselmann and Pokorny [23] consider the latent space in their formulation. More recently, formal methods such as temporal logic have been used to encode high-level constraints [15]. While these methods have shown empirical success in well-structured or open environments, they often struggle in cluttered settings where disconnected free spaces, narrow passages, or bottlenecks make it difficult for agents to reach informative states. The most closely related work is [24], which leverages RRT-based guidance to improve exploration. However, their method is fundamentally constrained by the feasibility of constructing valid planning trajectories, and lacks the extensibility needed to handle unseen obstacles or dynamic changes, both of which are addressed in our work.
3 Preliminary and Problem Formulation
@@
System Disturbance: We consider a continuous-time dynamical system $S$ with general unknown dynamics, whose evolution is described by: $\frac{ds}{dt}=\dot{s}=f(s,a,d)$ where $s\in\mathcal{S}\subseteq\mathbb{R}^{n}$ represents the state variables, $a\in\mathcal{A}\subseteq\mathbb{R}^{m}$ denotes the control inputs, and $d$ is the disturbance function.
@@
Figure 1: Comparison between previous method and graph-based method. We implemented both RRT (left) and RRG (middle) to solve the same goal-reaching task. As shown, the gray areas represent known obstacles, and the red cross indicates potential disturbances that may lead to failure transitions. The right figure shows a 3D navigation environment where graph-based planning generates feasible paths (white lines) through complex obstacle configurations. The translucent walls represent static barriers in the 3D space, demonstrating the scalability of our approach to high-dimensional planning problems.
@@
We focus on finding a feasible solution in challenging environments. While standard MDP theory guarantees the existence of an optimal policy under mild conditions, learning such a policy via model-free RL remains difficult in sparse-reward or long-horizon settings. These challenges are further exacerbated in practice by the clutter of the environment. Our work specifically addresses these issues by enhancing exploration through a graph-based guidance framework.
@@
We compared our approach with methods that use motion planning-generated paths to guide the training of reinforcement learning agents [24] (see Figure 1). To illustrate their limitations, we deliberately introduced an infeasibility (indicated by a red cross) along the planned geometric path to the goal. The tree-based method (left) fails to manage such disturbances, as it relies on a single, fixed trajectory without considering alternative routes, resulting in failure. Consequently, using waypoints from this path as guidance also fails. In contrast, our approach employs a graph that covers the entire workspace (middle), providing more robust guidance.
4 Overcoming Exploration
In Section 4, we introduce the graph-based reward and our design enabling arbitrary starting points.
4.1 Geometric Graph-based Motion planning Methods
@@
Steering Function: Given two states $s$ and $s^{\prime}$, the function $\text{Steer}(s,s^{\prime},\eta)$ returns a new state $s_{\text{new}}$ such that $s_{\text{new}}$ lies on the path connecting $s$ to $s^{\prime}$ and the distance $\|s-s_{\text{new}}\|_{2}\leq\eta$. The step size $\eta$ is a user-defined parameter that controls how far the path extends in a single step.
@@
Sampling and connection: The sampling function generates a new state $s_{\text{new}}$ in the free space, and the connection function attempts to connect $s_{\text{new}}$ to its nearest neighbors $\mathcal{N}(s_{\text{new}})$ using a geometric trajectory. For RRG, if a connection is successful, the edge is added to the graph, forming a cycle when possible.
@@
Given the graph $G=(V,E)$ constructed by RRG, if a node $x\in V$ lies within the goal region $\mathcal{S}_{G}$, we compute the optimal state trajectory $\mathbf{x}^{*}$ as a sequence of geometric states: $\mathbf{x}^{*}=\{x_{0},x_{1},\ldots,x_{d}\},$ where $x_{i}\in V\subset\mathcal{S}$ denotes nodes and $x_{d}\in\mathcal{S}_{G}$. In most planning pipelines, Dijkstra’s algorithm [25] is employed to extract the shortest waypoint sequence from the constructed graph. Beyond sampling-based graphs, grid-based graphs constitute another practical instantiation of our framework. In such cases, connectivity is predefined, and shortest paths are typically computed using algorithms such as A*. Additional details are provided in Appendix A.
4.2 Motion Planning Guided Reward
@@
After employing a graph-based motion planning method, we obtained a set of waypoints that span the entire graph. Based on this, we use an optimal path algorithm to compute the distance from each node to the goal region. We denote the optimal trajectory by $\mathbf{x^{*}}$, and the optimal is defined according to the length of trajectories. $s|\mathbf{x^{*}}$ denotes a state of path $\mathbf{x^{*}}$. We define the distance function $Dist:\mathcal{S}\times\mathcal{S}\times\mathcal{X}\rightarrow[0,\infty]$ to represent the Euclidean distance between two states along a given path, and define the cost function $Cost:\mathcal{S}\times\mathcal{X}\rightarrow[0,\infty]$ to represent the distance between one state and the destination $s_{d}$. Hence, the distance between two states along the optimal path can be expressed as $Dist(s_{1},s_{2}|\mathbf{x^{*}})=|Cost(s_{1}|\mathbf{x^{*}})-Cost(s_{2}|\mathbf{x^{*}})|$. Consequently, the graph $G$ is now extended to $G=(V,E,Cost)$, where we slightly abuse the notation, denoting $Cost$ as distance between the current node and the goal.
@@
The distance along the path enables us to accurately quantify the actual path length, making it more applicable in cluttered environments compared to the Euclidean distance. Based on this distance, we design the motion-planning reward scheme to guide the RL agent. Given the difficulty of reaching an exact state in continuous space, we define the norm r-ball to loosen the criteria for determining whether the agent has reached specific states. Formally, it is defined by $Ball_{r}(s)=\left\{s^{\prime}\in\mathcal{S}\mid\|s-s^{\prime}\|_{2}\leq r\right\}$, where $r$ is the radius. The selection of the r-ball radius can influence the performance of the algorithm. For simplicity, we use the default $r=\frac{\eta}{2}$ based on the steering function in all the following sections. We then define a progression function $D:\mathcal{S}\rightarrow[0,\infty]$ to determine whether the agent has moved toward the goal region and got closer to it. Formally,
@@
We adopt the definition of history from [26] with a slight modification $h_{t}=\{s_{1},a_{1},\dots,s_{t-1},a_{t-1}\}$ instead of $\{s_{1},a_{1},\dots,s_{t-1},a_{t-1},s_{t}\}$ , thereby the $D_{min}$ over a state-action sequence $\tau_{t}=h_{t}\bigcup\{s_{t},a_{t}\}$ is defined as $D_{min}(\mathbf{\tau})=\min_{s\in\mathbf{s}}{D(s)}$, where $\mathbf{s}$ denotes the state sequence in trajectory $\tau_{t}$. Ideally, we would check the change of $D_{min}$ at each step to determine whether the agent has moved closer to the goal region, i.e., check if $D_{min}(\tau_{t})<D_{min}(\tau_{t-1})$, and assign a positive reward $R_{+}$ accordingly. However, this design introduces a non-Markovian reward issue. In other words, $r(h_{t},s_{t},a_{t})\neq r(h^{\prime}_{t},s_{t},a_{t})$ could happen with different history $h_{t}$ and $h^{\prime}_{t}$. To address this problem, we denote $\mathcal{S}^{\times}=\{(s,D_{min})\mid s\in\mathcal{S},\,D_{min}\in\mathbb{R}\}$, so that $D_{min}$ can be directly retrieved from the augmented state. Correspondingly, the augmented reward function is defined as $\tilde{r}:\mathcal{S}^{\times}\times\mathcal{A}\times\mathcal{S}^{\times}\to\mathbb{R}$,
@@
This result ensures that our planning-augmented reward remains compatible with standard RL algorithms. In the remainder of this paper, we operate in the augmented state space $\mathcal{S}^{\times}$ to maintain this property. The detailed proof is presented in Appendix B.
@@
This proposition guarantees that the auxiliary rewards accelerate convergence without altering the original task objective. The proof is presented in Appendix C. The learning objective aims to find the optimal policy by optimizing the parameters $\theta$:
@@
4.3 Infeasible Paths Analysis
Recalling Sec. 3, our method specifically addresses the challenge of unexpected disturbances in the simulation, which often render paths infeasible and cause tracking errors while training. In contrast to tree-based methods, which focus on planning a single path without accounting for its robustness or feasibility, our approach leverages a graph-based motion planning strategy.
We highlight that the robustness of our method stems from the redundancy introduced by graph-based planning. Rather than following a single predetermined trajectory, the agent can flexibly choose among multiple feasible alternatives toward the goal. This redundancy is particularly beneficial in the presence of disturbances or partial map mismatches, where some paths may become infeasible during execution. As the graph becomes denser—either through increased sampling resolution or stronger connectivity—the number of distinct paths between the start and goal increases. Prior work in random geometric graphs shows that the connectivity and availability of alternative paths grow rapidly with the number of nodes and connection radius [29]. This property greatly enhances the likelihood that at least one safe and feasible path exists, improving resilience against both local failures and global perturbations.
Figure 2: Left: Maze layout. Right: Diverse goal-reaching paths from nearby starting points.